Company currently buys a part from a supplier for $13.55 per unit but is considering making the part itself next year. This year, they purchased 3,000 units of this part. Estimated costs to make the part are:

Per-Unit Total     
Direct materials $2.60    $7,800   
Direct labor 4.75    14,250   
Variable overhead 4.10    12,300   
Fixed overhead 4.50    13,500   
Total $15.95    $47,850   


Of the estimated fixed overhead, $5,130 are common costs that would be allocated to the part; the rest would be additional fixed overhead costs. X Company currently rents unused factory space to a tenant for $2,900; it will have to use this space to make the part.

2. If X Company will need 3,000 units again next year and continues to buy the part instead of making it, it will save

3. X Company is somewhat uncertain how many units of the part it will need next year. How many units would make the company indifferent between continuing to buy the part and making it?
Answer #1

2) Differential analysis

Make Buy
Direct material 7800
Direct labor 14250
variable manufacturing overhead 12300
Fixed manufacturing overhead (13500-5130) 8370
Opportunity Cost 2900
Purchase cost (3000*13.55) 40650
Total relevant cost 45620 40650

it will save = 45620-40650 = $4970

3) relevant cost of manufacturing = relevant cost of buying

11.45X+8370+2900 = 13.55X

-2.10X = -11270

X(Indifferent point) = 11270/2.1 = 5367 Units